OpenStudy (anonymous):

8.214285714285714 in to a fraction

OpenStudy (alexandervonhumboldt2):

16 digits after decimal means 10000000000000000 thus we have \[\huge\bf {8\frac{ 214285714285714 }{ 10000000000000000 }=simplify}\]

OpenStudy (mathstudent55):

Is it repeating ...285714..., or does it end just like you have it?

OpenStudy (anonymous):

I need it as a whole number please if possible

OpenStudy (alexandervonhumboldt2):

whole number? it is imposbbile

OpenStudy (anonymous):

oh

OpenStudy (alexandervonhumboldt2):

we can round, but it will not be an eaxct number

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

\[1/7 = 0.\dot14285\dot7\\ 1/70 = 0.0\dot14285\dot7\\[2ex] 8.2\dot14285\dot7 \\ \quad= 8.2+0.0\dot14285\dot7\\ \quad = 8+0.2+1/70\\ \quad = 560/70+14/70+1/70 \\ \quad = 575/70\\ \quad = 115/14\]